The wavelength of light emitted from a traffic light having a frequency of 5.82x1014 Hz is nm. 25.8 51.5 515 17.2 50.0 Submit Previous AnswersRequest Answer X Incorrect; Try Again; 5 attempts remaining
Answer #1

we will solve it by the equation Lembda= speed of light/frequency

here speed of light = 3*10^8 m/s and frequency= 5.82*10^14 hz and Lembda= wavelength

using the values in the formula

Lembda = 3*10^8/5.82*10^14= 5.15*10^-7 = 515*10^-9 m =515 nm